# A Survey on Transformer Compression

### 05 Feb 2024 v1

@@ 4.1.1 Overview of Pruning
Neural network pruning has long been recognized as an effective method for shrinking model size and speeding up model inference. The taxonomy of pruning methods can be quite complicated, including the sequential order of pruning and model training, structure specification, and the way to determine the pruned parameters [133]. Within the scope of this review, however, limiting the source model as a pre-trained large transformer for natural language processing [134, 4] or visual recognition [12, 26, 135, 136] puts forward several specific categories of techniques to be addressed (Fig. 5).
As pre-training accounts for most of the model performance for downstream tasks, pruning after training (PAT) [137] becomes a major choice. At a high level, the whole process consists of pre-training, pruning, and retraining performance recovery. However, the multi-task nature of pre-training and huge computational in the retraining phase also lead to critical issues to be addressed for pruning large transformer models, as detailed later.
On the other hand, the structure specification of pruning separates the techniques into unstructured and structured pruning. The former conducts pruning at the finest-grained level [138, 139], i.e., weight-wise pruning, which follows the optimization problem below,
@@
While the parameter size or memory usage are generally reduced in unstructured pruning, latency speed-up can not be guaranteed since the resulting model shape tends to be irregular, and acceleration requires specific hardware design. In contrast, specifying the eliminated structure as an entire layer, head, or other network unit results in structural pruning [140], which can generally shorten the latency on standard hardware. When the specification is a pattern suitable for speeding-up (e.g., a specific ratio of non-zero mask), the pruning is called semi-structural pruning [141].
Besides network parameters, reducing input size is also an appealing direction to explore. With the emphasis on removing redundant information between tokens, as both discovered in vision [33] and language domain [142] transformers, this type of algorithm also achieves decreased computation FLOPs with unpruned network parameter.
@@ 4.1.2 Pruning for Transformer-Based Large Language Models
Pruning granularity. In response to the considerable parameter increase in the LLM era, early attempts of pruning arise with the focus on unstructured and semi-structured pruning [143, 144, 145]. SparesGPT, for instance, conducts unstructured and semi-structured (mainly 2:4) pruning starting from OPT-175B [146] and BLOOM-176B [147] and achieve 50-60% sparsity with moderate perplexity increase and performance decrease on downstream datasets. This work demonstrates the feasibility of low-resource pruning in $>$100 B models with moderate sparsity.
@@
Another line of work mainly focuses on the structured pruning method to achieve more significant hardware speed-up. Language structural pruning has been conducted at multiple granularity. For instance, [148] discusses the redundancy of attention head, [149] proposes a task-specific extraction of pre-trained sub-networks, [150] focuses on FFN layers, and Block Pruning [151] is introduced to prune MHA and FFN parameters separately. In comparison, CoFi [152] proposes a more comprehensive coarse- and fine-grained pruning method that learns to generate masks at different granularity, i.e., FFN layers, FFN intermediate dimensions, MHA layers, Attention heads, and hidden dimensions to encourage the optimization in high sparsity domain. In this way, the pruning of a specific parameter can be determined at different granularity in an end-to-end learnable manner. When pruning the general domain pre-trained BERT model to typical downstream datasets, CoFi achieves more than 10×10$\times$10 × speed-ups with a limited accuracy gap. Later, this method was extended to a larger decoder-based LLM with a 7B parameter size [30]. With a pre-defined model shape, the pruned model shows better instruction tuning performance than models of similar size but trained from scratch. Also, specifying the model configuration with uniform size avoids irregularities in model shape and further increases the inference throughput. In contrast to pre-defining the model shape, [29] proposed first to identify groups of coupled structures considering the interdependency of parameters in FFN, MHA, and channel-wise groups and conduct the pruning process by group. Experiments at a relatively low compression rate (20%) demonstrate the different effectiveness of these grouping strategies.
@@
Pruning criteria. To determine the structure to be pruned, various metrics have been explored in the context of LLM pruning. For instance,  [143] follows [153] to find an optimal mask that can be used for weight reconstruction and uses Hessian Matrix (i.e., the second order indicator of loss change concerning the change of this parameter) to indicate the parameter importance, which is experimentally shown better than a magnitude criteria [154] and much faster than previous method AdaPrune [155]. In comparison, as noted by LLM Pruner [29], the first-order term should also be considered due to the shift of data used in the pruning process compared to the original language model training. Also, it proposes considering the weight importance at different levels and explores different ways of combining the importance information. Besides,  [144] also explores the combination of magnitude and norm of the input activations as the pruning criteria, which lowers the computational cost used for Hessian computation.
@@
Learning to prune. Besides determinating parameter importance with designed metrics, one can also perform extra training, commonly with regularization on sparsity [156, 157]. In particular, the success of LLM as a multi-task language processor promotes the development of a pruning method that preserves this multi-task nature, as opposed to the task-specific trails that are more feasible in the traditional pretrain-finetune paradigm [152]. In response to this motivation, recent works such as SIMPLE [158], LLM Pruner [29] and Sheared LLama [30] incorporate the causal language modeling loss in the pruning objective. However, this “general pruning” strategy raises another issue of data selection in the pruning process, as the domain gap between pre-training and pruning may hinder the importance of estimation/performance recovery. As detailed in [30], a dynamic batch loading strategy is applied to enhance adaptively input the data batch from a proper domain to balance the model performance as a multi-task learner.
@@
Computation cost for parameter pruning. Considering the huge parameter size, post-pruning retraining and parameter importance estimation may both generate significant computation burden for large language models. In an early attempt, [143] considers a one-shot strategy, i.e., only limited data are used to calibrate the Hessian Matrix in importance estimation, and no retraining was employed. From another perspective, LoRAPrune [159] shows that introducing parameter-efficient tuning such as Low-Rank Adaption can also decrease the computation cost in retraining while compensating for the performance degradation.
@@
Context and token pruning. Different from the case in the computer vision domain, language sequence can be very long, and long-range reference or induction also calls for the expansion of input tokens (i.e., context) to prompt better language understanding. Also, since the inference cost of LLM is in quadratic complexity with respect to sequence length, this expansion greatly hinders the efficient inference of language models. To tackle these problems, methods have been proposed since BERT era to prune redundancy in context or context attention calculation. For context pruning,  [160, 142] proposes to prune tokens across the transformer layers progressively and improves the throughput by several folds on GLUE benchmark with less than 1% accuracy loss. Among the second line of research, various sparse attention techniques have been proposed, including local attention restricted to nearby tokens [161, 162, 163], introducing global tokens [164, 162], content-based token grouping [37, 163], and attention back tracking [165]. With the advance of the modeling capacity of LLM, the context input for LLM is also growing readily. To provide a more adaptive attention selection, [31] proposes a sparse sigmoid-based selection scheme. Also, erasing tokens from the key-value cache makes this method more hardware-friendly in a decoder-based language model. While the experiments are mainly conducted on modest size GPT-2 model [3], the empirical results that up to 80% of the context can be successfully pruned with negligible degradation in perplexity indicates an appealing potential to explore along this direction.
@@ 4.1.3 Pruning for Transformer-Based Vision Model
  • ${}^{\dagger}$ Data is from 2:4 partial sparsity setting with GPU speeding up [143].
@@
  • ${}^{\star}$ Compression rate is about context tokens instead of parameters.
@@
Token and feature pruning. The early exploration of pruning for vision transformer pruning focuses on reducing information redundancy in feature space. In a preliminary study, [171] proposes learning the most informative feature dimensions and neglecting the remaining dimensions in the inference time. This is achieved by learning a diagonal real-value feature mask during pre-training with a sparsity regularizer and obtaining the hard feature mask using a defined threshold. From a different perspective, given the fact that token features in late ViT layers are quite similar in feature embedding, patch slimming [33] method is proposed to prune the patches from the output layer to the input layer. In particular, for each layer, it evaluates the impact of a patch on the final output feature and removes less important ones. Besides determining the patch importance by training dataset statistics, token sparsification can also be achieved by inserting a light-weight prediction network to indicate the importance of the input patches [33, 166, 172, 173, 167]. Note that since the self-attention can accept the token sequence regardless of its length, unstructured pruning of input tokens can be easily implemented in a hardware-friend manner for these token pruning methods. Furthermore, similar techniques are adopted in other transformer-based vision models such as DETR [174] to promote focus on object tokens [175, 176].
@@
Structured pruning. Apart from the redundancy in feature dimension and tokens, the heavy burden of matrix computation is another important axis to improve model efficiency. In an early work, [167] proposes to selectively optimize for sparsity by obtaining unstructured subnetworks and structured pruning for attention heads, besides importance-based token selection. [169] develops a structural pruning method to reduce the network’s width by pruning linear projection matrices as well as attention heads, together with depth pruning with inserted early-exit classifiers at each layer. Later on, more works emerged to explore the collaborative effect of parameter pruning across different transformer modules. In [177], a Hessian-based criteria is proposed to evaluate the parameter importance globally over learnable parameters. Together with the parameter redistribution strategy that promotes shorter latency, this method can also introduce novel efficient ViT structures named NViT. With an emphasis on speeding up continuous structure search, [32] performs a single-shot architecture search for all the ViT components with significantly shorter GPU hours needed. Also, by incorporating the interactions between the components, [168] manages to introduce adaptively the pruning ratio for each component, adding another degree of freedom in pruning. More recently, [34] proposes to use class labels to optimize the fully differentiable pruning operation at each layer.
@@ 4.1.4 Discussion
Pruning serves as a fundamental way to reduce the computation burden of the pre-trained transformer. Traditionally, this is achieved in an orthogonal manner by pruning input features or network parameters. In particular, we notice that different pruning methods can biasedly benefit particular application purposes. For large language models, structured network pruning fits more in the setting when a modest length (e.g., several thousand tokens) is used as input to speed up inference, as the defined hardware-friendly model shape has been proven to realize a good trade-off between acceleration and perplexity decrease [30]. In comparison, context pruning or sparse attention methods deserve more attention when context length gets longer, and the ceiling of pruning ratio for context for LLMs (e.g., 7B or more) has not been determined in a detailed manner, which is an interesting direction to explore. For vision transformers, the development of pruning methods has systematically tackled each angle of model design, making the current pruning process more flexible and closer to a continuous structural search task, which can be an excellent reference to the development of language domains. The training cost is another critical axis to consider in the pruning of large models, especially when incremental LLM pre-training is considered a component of the pruning process. Therefore, developing training-efficient pruning methods is attracting more attention. And this unmet algorithm need can be fulfilled by more accurately estimating the parameter (or block) sensitivity by limited data or more detailed exploration of the information stored hidden features during training/inference.
@@ 4.2 Efficient Architecture
Transformer [11] has become the de facto backbone architecture to develop various large vision and language models (LVMs and LLMs), making it possible to scale to billions of parameters. In this section, we review the efficient architecture design of large transformer-based models in both vision and language domains, i.e., efficient design on mainstream architecture, attention mechanism and feed forward network.

### 07 Apr 2024 v2

@@ 4.1.1 Overview of Pruning
Neural network pruning has long been recognized as an effective method for making models more compact and speeding up model inference. The taxonomy of pruning methods can be quite complicated, including the sequential order of pruning and model training, structure specification, and the way to determine the pruned parameters [134].This study, however, limits the scope of the source model as a pre-trained large Transformer for natural language processing [135, 4] and for visual recognition [12, 26, 136], raising several specific categories of techniques to be addressed (Fig. 5).
As pre-training accounts for most of the model performance for downstream tasks, pruning after training (PAT) [137] becomes a major choice. From a high-level perspective, the whole process consists of pre-training, pruning, and retraining performance recovery. However, the multi-task nature of pre-training and huge computational in the retraining phase also lead to critical issues to be addressed for pruning large Transformer models, as detailed later.
The methods available for pruning are generally split into two categories according to the structure specification of pruning: unstructured pruning and structured pruning. The former conducts pruning at the finest-grained level [138, 139], i.e., weight-wise pruning, which follows the optimization problem below,
@@
Although unstructured pruning generally leads to a reduction in the parameter size or memory usage, it cannot guarantee latency speedups because the resulting model shape tends to be irregular, and such speedups requires specific hardware design. In contrast, specifying the eliminated structure as an entire layer, head, or other network unit results in structural pruning [140], which can generally shorten the latency on standard hardware. When the specification is a pattern suitable for speeding-up (e.g., a specific ratio of non-zero mask), the pruning is called semi-structural pruning [141].
In addition to network parameters, reducing the input size is also an appealing direction to explore. With the emphasis on removing redundant information between tokens, such reduction also achieves lower computation FLOPs with unpruned network parameters, as discovered both in computer vision [33] and language domain [142] Transformers, .
@@ 4.1.2 Pruning for Transformer-Based Large Language Models
Pruning granularity. In response to the considerable parameter increase in the LLM era, there have been early attempts of pruning arise with the focus on unstructured and semi-structured pruning [143, 144, 145]. SparesGPT, for instance, conducts unstructured and semi-structured (mainly 2:4) pruning starting from OPT-175B [146] and BLOOM-176B [147] and achieve 50-60% sparsity with moderate perplexity increase and performance decrease on downstream datasets. This work demonstrates the feasibility of low-resource pruning in $>$100 B models with moderate sparsity.
@@
Another line of work mainly focuses on the structured pruning method to achieve a more significant speed-up. Language structural pruning has been conducted at multiple granularities. For instance, Michel et al. [148] discuss the redundancy of attention head, Fan et al. [149] propose a task-specific extraction of pre-trained sub-networks, Santacroce et al. [150] focus on FFN layers, and Block Pruning [151] is introduced to prune MHA and FFN parameters separately. In comparison, CoFi [152] proposes a more comprehensive coarse- and fine-grained pruning method that learns to generate masks at different granularity, i.e., FFN layers, FFN intermediate dimensions, MHA layers, Attention heads, and hidden dimensions to encourage the optimization in the high sparsity domain. In this way, the pruning of a specific parameter can be determined at different granularity in an end-to-end learnable manner. When pruning the general domain pre-trained BERT model to typical downstream datasets, CoFi achieves more than 10×10$\times$10 × speed-ups with a limited accuracy gap. Later, this method was extended to a larger decoder-based LLM with a 7B parameter size [30]. With a pre-defined model shape, the pruned model shows better instruction tuning performance than models of similar size but trained from scratch. Also, specifying the model configuration with uniform size avoids irregularities in model shape and further increases the inference throughput. In contrast to pre-defining the model shape, Ma et al. [29] propose first to identify groups of coupled structures considering the interdependency of parameters in FFN, MHA, and channel-wise groups and conduct the pruning process by group. Experiments at a relatively low compression rate (20%) demonstrate the varying effectiveness of these grouping strategies.
@@
Pruning criteria. To determine the structure to be pruned, various metrics have been explored in the context of LLM pruning. For instance, Frantar et al. [143] follows [153] to find an optimal mask that can be used for weight reconstruction and uses Hessian Matrix (i.e., the second order indicator of loss change concerning the change of each parameter) to indicate the parameter importance, which is experimentally shown better than a magnitude criteria [154] and much faster than previous method AdaPrune [155]. In comparison, as noted by LLM Pruner [29], the first-order term should also be considered due to the shift of data used in the pruning process compared to the original language model training. Also, it proposes considering the weight importance at different levels and explores different ways of combining the importance information. Besides, Sun et al. [144] also explores the combination of magnitude and norm of the input activations as the pruning criteria, which lowers the computational cost used for Hessian computation.
@@
Learning to prune. In addition to determining the importance of parameters using, one can also perform extra training, commonly with regularization on sparsity [156, 157]. In particular, the success of LLM as a multi-task language processor promotes the development of a pruning method that preserves this multi-task nature, as opposed to the task-specific trails that are more feasible in the traditional pretrain-finetune paradigm [152]. In response to this motivation, recent works such as SIMPLE [158], LLM Pruner [29] and Sheared LLama [30] incorporate the causal language modeling loss in the pruning objective. However, this “general pruning” strategy raises another issue of data selection in the pruning process, as the domain gap between pre-training and pruning may hinder the importance of estimation/performance recovery. As detailed in [30], a dynamic batch loading strategy is applied to adaptively input the data batch from a proper domain to balance the model performance as a multi-task learner.
@@
Computation cost for parameter pruning. Considering the huge numbers of parameters, post-pruning retraining and parameter importance estimation may both generate a significant computational burden for large language models. In an early attempt, Frantar et al. [143] consider a one-shot strategy, i.e., only limited data are used to calibrate the Hessian Matrix in importance estimation, and no retraining was employed. From another perspective, LoRAPrune [159] shows that introducing parameter-efficient tuning such as Low-Rank Adaption can also decrease the computation cost in retraining while compensating for the performance degradation.
@@
Context and token pruning. Different from the case in the computer vision domain, language sequence can be very long, and long-range reference or induction also calls for the expansion of input tokens (i.e., context) to prompt better language understanding. Also, since the inference cost of LLM is in quadratic complexity with respect to sequence length, this expansion greatly hinders the efficient inference of language models. To tackle these problems, methods have been proposed since BERT era to prune redundancy in context or context attention calculation. For context pruning, Kim et al. [160] propose to prune tokens across the Transformer layers progressively and improves the throughput by several folds on GLUE benchmark with less than 1% accuracy loss. Among the second line of research, various sparse attention techniques have been proposed, including local attention restricted to nearby tokens [161, 162, 163], introducing global tokens [164, 162], content-based token grouping [37, 163], and attention back tracking [165]. With the advance of the modeling capacity of LLM, the context input for LLM is also growing readily. To provide a more adaptive attention selection, Anagnostidis et al. [31] proposes a sparse sigmoid-based selection scheme. Also, erasing tokens from the key-value cache makes this method more hardware-friendly in a decoder-based language model. While the experiments are mainly conducted on modest size GPT-2 model [3], the empirical results that up to 80% of the context can be successfully pruned with negligible degradation in perplexity indicates an appealing potential to explore along this direction.
@@ 4.1.3 Pruning for Transformer-Based Vision Model
  • ${}^{\dagger}$ Data is from 2:4 partial sparsity setting with GPU speedup [143].
@@
  • ${}^{\star}$ Compression rate refers to context tokens instead of parameters.
@@
Token and feature pruning. The early exploration of pruning for vision Transformer pruning focuses on reducing information redundancy in feature space. In a preliminary study, Zhu et al. [171] propose learning the most informative feature dimensions and neglecting the remaining dimensions in the inference time. This is achieved by learning a diagonal real-value feature mask during pre-training with a sparsity regularizer and obtaining the hard feature mask using a defined threshold. From a different perspective, given the fact that token features in late ViT layers are quite similar in feature embedding, patch slimming [33] method is proposed to prune the patches from the output layer to the input layer. In particular, for each layer, it evaluates the impact of a patch on the final output feature and removes less important ones. Besides determining the patch importance by training dataset statistics, token sparsification can also be achieved by inserting a light-weight prediction network to indicate the importance of the input patches [33, 166, 172, 173, 167]. Note that since the self-attention can accept the token sequence regardless of its length, unstructured pruning of input tokens can be easily implemented in a hardware-friend manner for these token pruning methods. Furthermore, similar techniques are adopted in other Transformer-based vision models such as DETR [174] to promote focus on object tokens [175, 176].
@@
Structured pruning. Apart from the redundancy in feature dimension and tokens, the heavy burden of matrix computation is another important axis to improve model efficiency. In an early work, [167] proposes to selectively optimize for sparsity by obtaining unstructured subnetworks and structured pruning for attention heads, besides importance-based token selection. [169] develops a structural pruning method to reduce the network’s width by pruning linear projection matrices as well as attention heads, together with depth pruning with inserted early-exit classifiers at each layer. Later on, more works emerged to explore the collaborative effect of parameter pruning across different Transformer modules. In [177], a Hessian-based criteria is proposed to evaluate the parameter importance globally over learnable parameters. Together with the parameter redistribution strategy that promotes shorter latency, this method can also introduce novel efficient ViT structures named NViT. With an emphasis on speeding up continuous structure search, Chavan et al. [32] perform a single-shot architecture search for all the ViT components with significantly fewer GPU hours needed. Also, by incorporating the interactions between the components, Zheng et al. [168] manage to introduce adaptively the pruning ratio for each component, adding another degree of freedom in pruning. More recently, Yu et al. [34] propose to use class labels to optimize the fully differentiable pruning operation at each layer.
@@ 4.1.4 Discussion
Pruning serves as a fundamental way to reduce the computation burden of the pre-trained Transformer. Traditionally, this is achieved in an orthogonal manner by pruning input features or network parameters. In particular, we notice that different pruning methods can biasedly benefit particular application purposes. For large language models, structured network pruning fits more in the setting when a modest length (e.g., several thousand tokens) is used as input to speed up inference, as the defined hardware-friendly model shape has been proven to realize a good trade-off between acceleration and perplexity decrease [30]. In comparison, context pruning or sparse attention methods deserve more attention when context length gets longer, and the ceiling of pruning ratio for context for LLMs (e.g., 7B or more) has not been determined in a detailed manner, which is an interesting direction to explore. For vision Transformers, the development of pruning methods has systematically tackled each angle of model design, making the current pruning process more flexible and closer to a continuous structural search task, which can be an excellent reference to the development of language domains. The training cost is another critical axis to consider in the pruning of large models, especially when incremental LLM pre-training is considered a component of the pruning process. Therefore, developing training-efficient pruning methods is attracting more attention. This requires the parameter (or block) sensitivity to be estimated more accurately using limited data or the information stored in hidden features during training/inference to be explored more in-depth.
@@ 4.2 Efficient Architecture
Transformer [11] has become the de facto backbone architecture to develop various large vision and language models (LVMs and LLMs), making it possible to scale to billions of parameters. In this section, we review the efficient architecture design of large Transformer-based models in both vision and language domains, i.e., efficient design on mainstream architecture, attention mechanism and feed forward network. Notably, here we also discuss the recently proposed Mamba [178], a general sequence model backbone that integrates selective state space models (SSMs) into a simplified end-to-end neural network architecture. In this
section, we consider it as an efficient design for subquadratic-time architectures.
4.2.1 Representative Transformer-Based Models
Table VII presents the model cards of several representative LLMs and LVMs with public details. Vanilla Transformer [11] is based on the encoder-decoder architecture [133]. In this architecture, the encoder stacks multi-head self-attention layers to encode the input sequence into latent representations, while the decoder performs cross-attention on these representations and generates target sequences in an autoregressive manner. BERT [91] is designed to pre-train deep bidirectional representations from unlabeled text by considering both left and right context in all layers. GPT-series models [4, 2] have effectively demonstrated the power of in-context learning using decoder-only architectures [147, 179, 180]. These decoder architectures incorporate unidirectional attention masks to ensure that each input token can only attend to previous tokens, including itself. The input and output tokens are processed in a similar manner within the decoder. In the following, we will provide separate reviews from the perspectives of CV and NLP on the current structural innovations in Transformer-based architectures.
4.2.2 Language Domain
The self-attention mechanism in conventional Transformer architectures often faces quadratic computational complexity. This poses a challenge for training and making inferences with long input sequences. To enhance efficiency, current innovations in the mainstream can be classified into three categories: optimizations targeting the attention mechanism, the direct replacement of attention with more efficient architectures, and enhancements focused on the FFN component.
Attention innovation. Various efficient Transformer variants have been proposed to reduce the computational complexity of the attention mechanism by incorporating structural priors on attention, such as sparsity [181, 182, 183, 184, 185, 186]. For example, Reformer [37] approximates full attention computation using locality-sensitive hashing, reducing the complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$. Other techniques include pooling-based compression [185], clustering methods [184, 187] that apply k-means clustering to learn dynamic sparse attention regions, and Longformer [188], which combines local windowed attention with task-motivated global attention. Chelba et al. [189] proposes truncating the target-side window based on an $N$-gram assumption, restricting the self-attention mechanism to use only the previous $N-1$ tokens. Additionally, locally banded sparse attention methods, such as Factorized Attention [183], have been adopted in models like GPT-3 [4].
Other methods have been proposed to reduce the computational complexity of the attention mechanism using linear attention. Performer [190] introduces fast attention via positive orthogonal random features, which is a scalable kernel method for approximating softmax attention kernels. This approach can efficiently model kernelizable attention mechanisms beyond softmax and comes with strong theoretical guarantees, including nearly unbiased estimation of the attention matrix, uniform convergence, and low estimation variance. Linear Transformer [191] reformulates self-attention as a linear dot-product of kernel feature maps. It uses kernels $\frac{\phi(q_{i})\phi(k_{j})}{\sum_{n=1}^{|x|}\phi(q_{i})\phi(k_{n})}$ to replace the softmax function and takes advantage of the associativity property of matrix products to reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$.
Moreover, there have been studies [35, 192, 193] exploring the concept of multi-query attention, where different heads share the same linear transformation matrices for the keys and values, as proposed by Shazeer et al.[193]. This approach offers significant computational savings with only a minimal impact on model quality. To further strike a balance between multi-query attention and multi-head attention, GQA et al. [194] introduces the concept of grouped attention heads, where heads within the same group share identical transformation matrices.